# On Viscosity Solutions of Hamilton-Jacobi Equations in the Wasserstein space and the Vanishing Viscosity Limit

### v1

@@ 2.6.1. First order HJE
In order to deal with the first-order equation, the author of [BSOT] considers the following relaxed Hamiltonian $\mathcal{H}\colon\mathcal{P}_{2}(\mathbb{R}^{2d})\to\mathbb{R}$
(2.16) ${\mathcal{H}}(\gamma)=\int_{\mathbb{R}^{2d}}H(x,z,{\pi_{1}}_{\#}\gamma)d\gamma(x,z).$
This relaxation is reminiscent of the Kantorovich relaxation in the classical optimal transport theory and is the main ingredient of the comparison principle in [BSOT, Theorem 2.11]. We stress that the relaxed Hamiltonian coincides with the usual one at elements induced by maps, and, by the growth condition on $H$,
$\mathcal{H}(\gamma)\leq C(1+\|\gamma\|^{2}_{(\pi_{1})_{\#}\gamma}+\mathcal{M}_{2}((\pi_{1})_{\#}\gamma)),\quad\forall\gamma\in\mathcal{P}_{2}(\mathbb{R}^{2d}).$
@@
We say that $U$ is a viscosity solution of (1.1) if it is a subsolution and a supersolution.
We say that $U:[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is a classical sub(super)solution if $U$ is $C^{1}$ and
$\partial_{t}U(t,\mu)+\int_{\mathbb{R}^{d}}H(x,\mu,\partial_{\mu}U)d\mu\leq(\geq)0,\quad\forall(t,\mu)\in[0,T)\times\mathcal{P}_{2}(\mathbb{R}^{d}).$
@@
We recall that the notion of viscosity solution introduced in [BSOT, Definition 2.8] is obtained from Definition 2.16 by replacing the differential $\partial$ with the larger set $\partial_{S,I}$. We emphasize that the notion of viscosity solution adopted in Definition 2.16 is weaker than the one considered in [BSOT]. Indeed, any subsolution (resp. supersolution) in the sense of [BSOT, Definition 2.8] is also a subsolution (resp. supersolution) in the sense of Definition 2.16.
@@
Moreover, the two notions can be compared by means of the comparison principle presented in [BSOT]. In particular, one can prove that they coincide when the underlying space is the flat torus $\mathbb{T}^{d}$, provided that existence holds for both notions of solution. This equivalence follows from Proposition 2.14. For an extension of this result to the whole space $\mathbb{R}^{d}$, one may additionally assume that both solutions are continuous with respect to the weak convergence of measures. This assumption guarantees the existence of a maximum in the doubling variables argument.
@@ 2.6.2. Semilinear HJE with idiosyncratic noise
We start by recalling, for the sake of completeness, the definition of Entropy functional and Fisher information. We refer the reader to [AGS] and [OTVillani] for more details.
The entropy functional $\mathcal{E}\colon\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is defined as
$\mathcal{E}(\mu)=\begin{cases}\int_{\mathbb{R}^{d}}\mu\log\mu dx,\quad&\text{if }\mu\in\mathcal{P}_{2,ac}(\mathbb{R}^{d})\\ \infty\quad&\text{otherwise}.\end{cases}$
Here, and throughout the paper, we adopt the usual abuse of notation by denoting by $\mu$ also the density of the measure $\mu$ with respect to the Lebesgue measure $\lambda_{d}$, whenever it exists. Having formally, $\partial_{\mu}\mathcal{E}(\mu)=\nabla\frac{\delta}{\delta\mu}\mathcal{E}(\mu)=\nabla\log\mu$ the Fisher information is defined by
$\mathcal{I}(\mu)=\|\partial_{\mu}\mathcal{E}(\mu)\|_{L^{2}_{\mu}(\mathbb{R}^{d})}^{2}=\int_{\mathbb{R}^{d}}|\nabla\log\mu|^{2}dx=\int_{\mathbb{R}^{d}}\frac{|\nabla\mu|^{2}}{\mu}dx.$
@@
for $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, see [DS, Equation  (2.5)].
For the definition of viscosity solution of (1.2), we follow [DS]. Note that the authors introduce an entropic penalization that can be absorbed, thanks to the semilinear structure, into the equation, up to a constant depending only on the magnitude of the penalization itself and the diffusion matrix.
Definition 2.18 (See [DS, Definition 2.10]).
@@
We say that $U$ is a viscosity solution of (1.2) if it is both a subsolution and a supersolution.
@@
In view of the computation done in [DS, Section  2.3], the constant $K$ is expected to blow up as the diffusion goes to zero. We will take care of this fact in the proof of the vanishing viscosity limit.
@@ 3. Hopf-Lax formula
The goal of this section is to establish an analogue of the finite-dimensional Hopf–Lax formula in the Wasserstein space. We show that this formula provides a representation of the viscosity solution to (1.1) in the sense of Definition 2.16. A first extension in this direction was obtained in [GANGBO2019119] for the quadratic case, where the specific structure of the Hamiltonian is essential due to its relation with a projection operator. Here, we generalize their result beyond the quadratic setting, allowing for a convex Hamiltonian H depending solely on the adjoint variable. Moreover, in contrast with the assumptions in [BSOT, Section 5], we do not require differentiability of the associated Lagrangian. In particular, the Hopf–Lax formula still yields a viscosity solution even when the regularity hypotheses ensuring existence in [BSOT] are not satisfied.
@@
The above formula is formally derived using the method of characteristics. Its power also lies in the fact that it provides an alternative approach to obtain properties such as finite speed propagation of characteristics, Lipschitz continuity of the solution, and the propagation of semiconcavity provided the datum $g(\cdot)$ is semiconcave. Remarkably, in case the Hamiltonian $H$ is strictly convex, the semiconcavity of the solution at positive time is ensured even in the case of a non semiconcave datum. Let us underline that the key ingredient to prove that the function $u$, given by Hopf-Lax formula, is the viscosity solution of the HJ equation is the semigroup structure given in the form of a Dynamic Programming Principle:
@@
The aforementioned properties are standards for finite dimensional Euclidean spaces, see [Evans, Section 3.3], and can be classically proven also for Hilbert spaces, see [lasry_remark_1986] and references therein.
In this section, we make the following assumptions:
@@
    $L(z)\geq c\|z\|^{2}-C\quad\forall\,z\in\mathbb{R}^{d}.$
We start by considering a relaxation of the Lagrangian
@@ 3.1. The relaxed Lagrangian as the Fenchel transform of the relaxed Hamiltonian
$\mathcal{L}\colon\mathcal{P}_{2}(\mathbb{R}^{2d})\to\mathbb{R}$, defined by
@@
Let $\mathcal{H}:\mathcal{P}_{2}(\mathbb{R}^{2d})\to\mathbb{R}$ be the relaxed Hamiltonian associated to $H(p)=\sup_{v\in\mathbb{R}^{d}}\big\{\big<-v,p\big>-L(v)\big\}$, i.e. $L^{*}=H$ defined in (2.16). Fix $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Then, the equalities
$\mathcal{H}(\gamma)$ $=\sup_{\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}),\theta\in\Gamma_{\mu}(\gamma,\xi)}\Big\{\int_{\mathbb{R}^{3d}}\big(\big<-v,p\big>-L(v)\big)d\theta(x,p,v)\Big\}$
$=\sup_{\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}),\theta\in\Gamma_{\mu}(\gamma,\xi)}\Big\{\big<-\xi,\gamma\big>_{\theta,\mu}-\mathcal{L}(\xi)\Big\}$
$=\sup_{\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})}\Big\{\big<-\xi,\gamma\big>_{\mu}-\mathcal{L}(\xi)\Big\}$
@@
Moreover, for $\theta\in\Gamma_{\mu}(\gamma,\xi)$
(3.4) $\mathcal{H}(\gamma)+\mathcal{L}(\xi)=\big<-\xi,\gamma\big>_{\theta,\mu}\Longleftrightarrow\mathrm{supp}(\theta)\subseteq\big\{(x,p,v)\in\mathbb{R}^{3d}:p\in-D^{-}L(v)\big\},$
where $D^{-}L$ denotes the subdifferential of $L$ in the classic convex sense, and the set of $\theta\in\Gamma_{\mu}(\gamma,\xi)$ satisfying (3.4) is not empty since it contains at least an element induced by a measurable map.
@@
Equivalently, $\mathcal{H}(\gamma)\geq\sup_{\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}),\theta\in\Gamma_{\mu}(\gamma,\xi)}\Big\{\int_{\mathbb{R}^{3d}}\big(\big<-v,p\big>-L(v)\big)d\theta(x,p,v)\Big\}$. The equality holds iff $\theta$ is concentrated on the multigraph $\text{Graph}({-D^{-}L})=\big\{(x,p,v)\in\mathbb{R}^{3d}:p\in-D^{-}L(v)\big\}$ by the characterization of optimality in the Fenchel transform. Note that $\text{Graph}({-D^{-}L})$ is closed, and it is not empty, since $L$ is convex and superlinear. Thereofore, for a fixed $\gamma\in\mathcal{P}_{2}(\mathbb{R}^{2d})$, we can find a Borel measurable function $f:\mathbb{R}^{2d}\to\mathbb{R}^{d}$ s.t.
$(x,f(x,v),v)\in\text{Graph}({-D^{-}L})\quad\gamma-(x,v)\,a.e.$
see Kuratowski-Ryll-Nardzewski Theorem [Bogachev2007, Thm 6.9.3]. In particular, $\theta:=(\text{Id}\times f\times\text{Id})_{\#}{\gamma}$ is s.t. $(\pi_{1},\pi_{2})_{\#}\theta=\gamma$ and it is concentrated on $\text{Graph}({-D^{-}L})$.

### v2

@@ 2.7.1. First order HJE
In order to deal with the first-order HJE, the author of [9] considers the following relaxed Hamiltonian $\mathcal{H}\colon\mathcal{P}_{2}(\mathbb{R}^{2d})\to\mathbb{R}$
(2.19) ${\mathcal{H}}(\gamma):=\int_{\mathbb{R}^{2d}}H(x,z,{\pi_{1}}_{\#}\gamma)d\gamma(x,z).$
This relaxation is reminiscent of the Kantorovich relaxation in the classical Optimal Transport theory and is the main ingredient of the comparison principle in [9, Theorem 2.11]. We stress that the relaxed Hamiltonian coincides with the usual one at elements induced by maps, and, by the growth condition on $H$,
(2.20) $|\mathcal{H}(\gamma)|\leq C(1+\|\gamma\|^{2}_{(\pi_{1})_{\#}\gamma}+\mathcal{M}_{2}((\pi_{1})_{\#}\gamma)),\quad\forall\gamma\in\mathcal{P}_{2}(\mathbb{R}^{2d}).$
@@
We say that $U$ is a viscosity solution of (1.1) if it is a viscosity subsolution and a viscosity supersolution.
We say that $U:[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is a classical sub(super)solution of (1.1) if $U$ is fully $C^{1}$ and
$\partial_{t}U(t,\mu)+\int_{\mathbb{R}^{d}}H(x,\partial_{\mu}U,\mu)d\mu(x)\leq(\geq)0,\quad\forall(t,\mu)\in[0,T)\times\mathcal{P}_{2}(\mathbb{R}^{d}).$
@@
We recall that the notion of viscosity solution introduced in [9, Definition 2.8] is obtained from Definition 2.19 by replacing the differential $\partial$ with the larger set $\partial_{S,I}$. We emphasize that the notion of viscosity solution adopted in Definition 2.19 is weaker than the one considered in [9]. Indeed, any subsolution (resp. supersolution) in the sense of [9, Definition 2.8] is also a subsolution (resp. supersolution) in the sense of Definition 2.19.
@@
Moreover, the two notions can be compared by means of the comparison principle introduced in [9]. In particular, one can show that they coincide when the underlying space is the flat torus $\mathbb{T}^{d}$, provided that existence holds for both notions of solution. This equivalence is a direct consequence of Proposition 2.15. An extension of this result to the whole space $\mathbb{R}^{d}$ can be obtained by additionally assuming that both solutions satisfy a weak continuity property. This assumption ensures the existence of a maximum in the doubling variables argument.
@@ 2.7.2. Semilinear HJE with idiosyncratic noise
We start by recalling, for the sake of completeness, the definition of Entropy functional and Fisher information. We refer the reader to [2] and [50] for more details.
The entropy functional $\mathcal{E}\colon\mathcal{P}_{2}(\mathbb{R}^{d})\to(-\infty,+\infty]$ is defined as
$\mathcal{E}(\mu):=\begin{cases}\int_{\mathbb{R}^{d}}\log\mu(x)d\mu(x),\quad&\text{if }\mu\in\mathcal{P}_{2,ac}(\mathbb{R}^{d})\\ +\infty\quad&\text{otherwise}.\end{cases}$
Here, and throughout the paper, we adopt the usual abuse of notation by denoting with $\mu$ also the density of the measure $\mu$ with respect to the Lebesgue measure $\lambda_{d}$, whenever it exists. Having formally, $\partial_{\mu}\mathcal{E}(\mu)=\nabla\mathcal{D}_{\mu}\mathcal{E}(\mu)=\nabla\log\mu(x)$ the Fisher information $\mathcal{I}\colon\mathcal{P}_{2}(\mathbb{R}^{d})\to[0,+\infty]$ is defined by
$\mathcal{I}(\mu):=\begin{cases}\int_{\mathbb{R}^{d}}|\nabla\log\mu(x)|^{2}d\mu(x)=\int_{\mathbb{R}^{d}}\frac{|\nabla\mu(x)|^{2}}{\mu(x)}dx\quad&\text{if }\mu\in\mathcal{P}_{2,ac}(\mathbb{R}^{d}),\quad\frac{|\nabla\mu(x)|^{2}}{\mu(x)}\in L^{1}(\mathbb{R}^{d})\\ +\infty\quad&\text{otherwise}.\end{cases}$
@@
for $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, see [30, Equation  (2.5)].
For the definition of viscosity solution of (1.2), we follow [30]. Note that the authors introduce an entropic penalization that, thanks to the semilinear structure, can be absorbed into the equation, up to a constant depending only on the magnitude of the penalization itself and the diffusion matrix.
Definition 2.21 (See [30, Definition 2.10]).
@@
We say that $U$ is a viscosity solution of (1.2) if it is both a viscosity subsolution and a viscosity supersolution.
We say that $U:[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is a classical sub(super)solution of (1.2) if $U$ is partially $C^{2}$ and
$\partial_{t}U(t,\mu)-\int_{\mathbb{R}^{d}}\mathrm{tr}\!\left[a(\mu,x)\,\nabla\partial_{\mu}U\right]d\mu(x)+\int_{\mathbb{R}^{d}}H(x,\partial_{\mu}U,\mu)d\mu(x)\leq(\geq)0,\quad\forall(t,\mu)\in[0,T)\times\mathcal{P}_{2}(\mathbb{R}^{d}).$
@@
In view of the computation done in [30, Section  2.3, Eq, (2.23)], the constant $K$ is expected to blow up as the diffusion goes to zero. We will take care of this fact in the proof of the vanishing viscosity limit.
@@ 3. Hopf-Lax formula
The goal of this section is to establish an analogue of the finite-dimensional Hopf–Lax formula in the Wasserstein space. We show that this formula provides a representation of the viscosity solution to (1.1) in the sense of Definition 2.19. A first extension in this direction was obtained in [36] for the quadratic case, where the specific structure of the Hamiltonian is essential due to its relation with a projection operator. Here, we generalize their result beyond the quadratic setting, allowing for a convex Hamiltonian $H$ depending solely on the adjoint variable. Moreover, in contrast with the assumptions in [9, Section 5], we do not require differentiability of the associated Lagrangian. In particular, the Hopf–Lax formula still yields a viscosity solution even when the regularity hypotheses ensuring existence in [9] are not satisfied.
@@
The above formula is formally derived using the method of characteristics. Its power also lies in the fact that it provides an alternative approach to obtain properties such as finite speed propagation of characteristics, Lipschitz continuity of the solution, and the propagation of semiconcavity provided the datum $g(\cdot)$ is semiconcave. Remarkably, in case the Hamiltonian $H$ is strictly convex, the semiconcavity of the solution at positive time is ensured even in the case of a non semiconcave datum. Let us underline that the key ingredient to prove that the function $u$, given by Hopf-Lax formula, is the viscosity solution of the HJE is the semigroup structure given in the form of a Dynamic Programming Principle:
@@
The aforementioned properties are standards for finite dimensional Euclidean spaces, see [32, Section 3.3], and can be classically proven also for Hilbert spaces, see [40] and references therein.
In this section, we make the following assumptions focusing on the Lagrangian, rather than on the Hamiltonian:
@@
    $L(z)\geq c|z|^{2}-C\quad\forall\,z\in\mathbb{R}^{d}.$
@@ 3.1. The relaxed Lagrangian as the Fenchel transform of the relaxed Hamiltonian
We start by considering a relaxation of the Lagrangian $\mathcal{L}\colon\mathcal{P}_{2}(\mathbb{R}^{2d})\to\mathbb{R}$, defined by
@@
Let $\mathcal{H}:\mathcal{P}_{2}(\mathbb{R}^{2d})\to\mathbb{R}$ be the relaxed Hamiltonian associated to $H(p)=\max_{v\in\mathbb{R}^{d}}\big\{\big<-v,p\big>-L(v)\big\}$, i.e. $L^{*}=H$, defined in (2.19). Fix $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Then, the equalities
$\mathcal{H}(\gamma)$ $=\max_{\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}),\theta\in\Gamma_{\mu}(\gamma,\xi)}\Big\{\int_{\mathbb{R}^{3d}}\big(\big<-v,p\big>-L(v)\big)d\theta(x,p,v)\Big\}$
$=\max_{\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}),\theta\in\Gamma_{\mu}(\gamma,\xi)}\Big\{\big<-\xi,\gamma\big>_{\theta,\mu}-\mathcal{L}(\xi)\Big\}$
(3.4) $=\max_{\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})}\Big\{\big<-\xi,\gamma\big>_{\mu}-\mathcal{L}(\xi)\Big\}$
@@
Moreover, for $\gamma\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$
(3.5) $\mathcal{H}(\gamma)+\mathcal{L}(\xi)=\big<-\xi,\gamma\big>_{\mu}\Longleftrightarrow\xi=(\pi_{1},-D_{p}H(\pi_{2}))_{\#}\gamma.$
In particular,
(3.6) $\mathcal{H}(\gamma)=\max_{\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})}\Big\{-\big<\xi,\gamma\big>_{\mu}-\mathcal{L}(\xi)\Big\}.$
If, in addition,
(3.7) $\mathcal{L}(\xi)\geq\mathcal{L}(\mathrm{P}(\xi)),\quad\forall\xi\in\mathcal{P}_{2}(\mathbb{R}^{2d}),$
then, for all $\gamma\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$,
(3.8) $\mathcal{H}(\gamma)=\max_{\xi\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})}\Big\{\langle-\xi,\gamma\rangle_{\mu}-\mathcal{L}(\xi)\Big\}=\max_{\xi\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})}\Big\{-\langle\xi,\gamma\rangle_{\mu}-\mathcal{L}(\xi)\Big\},$
In particular, any optimizer $\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})$ in (3.5), with $\gamma\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$, is s.t. $\xi=\mathrm{P}(\xi)$, i.e. $\xi\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ is the only maximizer in (3.8).
@@
Equivalently,
$\mathcal{H}(\gamma)\geq\sup_{\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}),\theta\in\Gamma_{\mu}(\gamma,\xi)}\Big\{\int_{\mathbb{R}^{3d}}\big(\big<-v,p\big>-L(v)\big)d\theta(x,p,v)\Big\}.$
The equality holds iff $\theta$ is concentrated in $\Gamma_{H}=\big\{(x,p,v)\in\mathbb{R}^{3d}:v=-D_{p}H(p)\big\}$, where we denoted by $D_{p}H$ the differential of $H=L^{*}$, that is $C^{1}$ because $L$ is strictly convex and superlinear. This implies that $\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d}),\theta\in\Gamma_{\mu}(\gamma,\xi)$ optimize the first equality iff $\xi=(\pi_{1},-D_{p}H(\pi_{2}))_{\#}\gamma$, and $\theta=(\pi_{1},\pi_{2},-D_{p}H(\pi_{2}))_{\#}\gamma$. This gives the assertion (3.5). In particular, we also derive that for such a $\xi$
$\left<-\gamma,\xi\right>_{\mu}$ $=\max_{\theta\in\Gamma_{\mu}(\gamma,\xi)}\left<-\gamma,\xi\right>_{\mu,\theta}\underbrace{=}_{\Gamma_{\mu}(\gamma,\xi)=\{(\pi_{1},\pi_{2},-D_{p}H(\pi_{2}))_{\#}\gamma\}}\int_{\mathbb{R}^{2d}}\left<p,D_{p}H(p)\right>d\gamma$
$=-\int_{\mathbb{R}^{2d}}\left<p,-D_{p}H(p)\right>d\gamma=-\left<\gamma,\xi\right>_{\mu},$
that gives (3.6).
Now, suppose the decreasing property (3.7). For $\gamma\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$, we have
$\mathcal{H}(\gamma)$ $=\max_{\xi\in\mathcal{P}_{2,\mu}(\mathbb{R}^{2d})}\left\{\left<-\gamma,\xi\right>_{\mu}-\mathcal{L}(\xi)\right\}$
[…]
$=\sup_{\xi\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})}\left\{\left<-\gamma,\xi\right>_{\mu}-\mathcal{L}(\xi)\right\}.$
Moreover, if $\xi$ is an optimizer of (3.1) then $\text{P}(\xi)$ is an optimizer of (3.1) :
[…]
$\underbrace{=}_{\gamma\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})}\left<\text{P}(\xi),-\gamma\right>_{\mu}\leq\mathcal{H}(\gamma)+\mathcal{L}(\text{P}(\xi)).$
This implies $\mathcal{H}(\gamma)+\mathcal{L}(\text{P}(\xi))=\left<\text{P}(\xi),-\gamma\right>_{\mu}$, i.e. $\text{P}(\xi)$ is an optimizer, and $\mathcal{L}(\text{P}(\xi))=\mathcal{L}(\xi)$. By the characterization of optimizers $\xi=\text{P}(\xi)=(\pi_{1},-D_{p}H(\pi_{2}))_{\#}\gamma$ and
$\mathcal{H}(\gamma)=\max_{\xi\in\mathrm{Tan}^{K}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})}\left\{\left<-\gamma,\xi\right>_{\mu}-\mathcal{L}(\xi)\right\}.$